OpenStudy (anonymous):

Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions and identify what each term and factor represents in terms of the cell phone plans.

OpenStudy (texaschic101):

Can you make up prices to do this problem ?

OpenStudy (anonymous):

yes you can :)

OpenStudy (texaschic101):

Ameri-mobile has low activation fee but high monthly payments activation fee $50....monthly payment $ 100 A(ameri-mobile) = 50 + 100m -- m being months Cell-u-later has high activation fee and low monthly payments activation fee $ 100 ...monthly payment $ 50 C (cell-u-later) = 100 + 50m As far as finding out how much each plan costs, sub in the number of months for m. do I need to do examples, lets say, if you had it for 7 months ? Or do you need that ?

OpenStudy (anonymous):

no, not really, I just need to identify the terms or something. They don't show a number of months or anything... so... yeeeeaaah... Lol

OpenStudy (texaschic101):

do you understand how I made the equations ? Do you understand the terms ?

OpenStudy (anonymous):

Yeah, the terms are the letters representing the companies right? I'm sorry I had a rough week and practically everything slipped my mind so somethings are jumbled up !

OpenStudy (texaschic101):

yes... company A (ameri-mobile) = 50(activation) + 100m (m being number of months) company C (cell-u-later = 100(activation) + 50m does that make sense ?

OpenStudy (texaschic101):

oh...and the 100 and the 50 are what the two plans charge per month

OpenStudy (anonymous):

yup! perfect sense. M= months C=Cell-U-Later A=Ameri-Mobile and the numbers are made up. but yeah, I think my teacher would understand it. Right?

OpenStudy (texaschic101):

I believe so :)

OpenStudy (anonymous):

Yay! Thanks soooooo muuuch! I really needed that to have a good answer! haha, thanks xx

OpenStudy (texaschic101):

glad to help :)